If a driver provides 3.07 g of ethanol (C2H6O) into the Breathalyzer, what is the concentration of acetic acid (C2H4O2) if we assume the final volume is 100. mL?
Answer #1

Moles of ethanol = mass / molar mass

= 3.07 / 46

= 0.0667 moles

In breathalyzer equation

3C2H5OH + 2K2Cr2O7 + 8H2SO4 = 2Cr2(SO4)3 + 3CH3COOH + 2K2SO4 + 11 H2O

so mole of C2H5OH = mole of CH3COOH = 0.0667

So Molarity = mole / volume ( in litre)

= 0.0667 / 0.1

= 0.667 M

If you satisfied with the solution please like it thankxx

If you have any query please comment